1) Calculate the pH of the solution that results from each of the following mixtures. A) 150.0 mL...

1)

Calculate the pH of the solution that results from each of the following mixtures.

A)

150.0 mL of 0.26 M  HF with 225.0 mL of 0.32 M  NaF

B)

165.0 mL of 0.12 M C2H5NH2 with 275.0 mL of 0.22 M C2H5NH3Cl

2)

A)

As a technician in a large pharmaceutical research firm, you need to produce 100. mLof a potassium dihydrogen phosphate buffer solution of pH = 6.91. The pKa of H2PO4− is 7.21.

You have the following supplies: 2.00 L of 1.00 M KH2PO4 stock solution, 1.50 L of 1.00 M K2HPO4 stock solution, and a carboy of pure distilled H2O.

How much 1.00 M KH2PO4 will you need to make this solution? (Assume additive volumes.)

B)

If the normal physiological concentration of HCO3− is 24 mM, what is the pH of blood if PCO2 drops to 31.0 mmHg ?

3)Which pair of concentrations results in the most effective buffer?

  

0.95 MHA; 0.05 M A−

0.10 MHA; 0.90 M A−

0.20 MHA; 0.20 M A−

0.70 MHA; 0.70 M A−
Answer #1

1) A) Moles of HF, a weak acid = 0.15 * 0.26 = 0.039 and moles of F-, its conjugate base = 0.225 * 0.32 = 0.072
Total volume of the mixture = 0.15 + 0.225 L = 0.375 L
Hence, New [HF] = 0.039/0.375 = 0.104 M and New [F-] = 0.072/0.375 = 0.192 M
pH of a weak acid buffer = pKa + log([HF]/[F-])
= -log(6.6*10-4) + log[(0.104)/(0.192)]   [Ka value of HF = 6.6 * 10-4]
= 2.91

B) Moles of C2H5NH2 = 0.165 * 0.12 = 0.019 and moles of C2H5NH3+ = 0.275 * 0.22 = 0.06
Total volume = 0.44 L
Thus, [C2H5NH2] = 0.019/0.44 = 0.043M and [C2H5NH3+] = 0.06/0.44 = 0.136M
pOH of a weak base buffer = pKb + log([C2H5NH3+]/[C2H5NH2] )
= -log (4.3*10-4) + log[(0.136)/(0.043)]
= 3.86.
  Thus, pH = 14-3.86 = 10.14

2.A) pH of the buffer = pKa + log[HPO4-2]/log[H2PO4-]
i.e. 6.91 = 7.21 + log[HPO4-2]/[H2PO4-] ; [HPO4-2]/[H2PO4-] = 0.5
Let the volume of KH2PO4 be x L. Thus, volume of K2HPO4= 0.25-x L.
Now,
moles of KH2PO4 = 2 and moles of K2HPO4 = 1.5
Thus, [HPO4-2]/[H2PO4-] = 0.5 = (1.5/0.25-x) / (2/x)
Solving for x, we obtain
x, i.e. volume of KH2PO4 = 0.1 L = 100 mL.